OpenStudy (anonymous):

please help. Calculate the consumers' surplus at the indicated unit price p' for the demand equation. q = 50 − 3p; p' = 13

OpenStudy (anonymous):

I am kinda confuse here about what you are trying to ask. So p=13 so u plug that into the equation which is q=50-3(13) which leads to q=50-39. then u can solve from there.

OpenStudy (anonymous):

It is consumer surplus I dont know how to solve it

OpenStudy (anonymous):

Dont you just plug the p into the equation though to find q?

OpenStudy (anonymous):

I think Im supposed to do something with this but I dont know \[\int\limits_{0}^{37/3} (q-50/-3) -13\]